Three charges are at the corners of an equilateral triangle with sides 9m. What is the potential energy of these charges.( q1 = q2 = 4×10(-4)C and q3 = -3×10(-4)C and K=9×109 )

1
Expert's answer
2021-04-26T18:42:16-0400

Potential Energy of "q_1" with respect to "q_2"


"U_{12}=K\dfrac{q_1q_2}{r}=9\times 10^{9}\times \dfrac{4\times 10^{-4}\times 4\times 10^{-4}}{9}=160J"


Potential Energy of "q_2" with respect to "q_3"


"U_{23}=K\dfrac{q_2q_3}{r}=9\times 10^{9}\times \dfrac{4\times 10^{-4}\times -3\times 10^{-4}}{9}=-120J"


Potential Energy of "q_3" with respect to "q_1"


"U_{13}=K\dfrac{q_1q_3}{r}=9\times 10^{9}\times \dfrac{4\times 10^{-4}\times -3\times 10^{-4}}{9}=-120J"


Net energy "U=U_{12}+U_{23}+U_{13}=160-120-120=-80J"